OpenStudy (anonymous):

how do i find the critical numbers for cube root of X+2 . I took the derivative of this and got (1)/[3(x+2)^(-2/3)]. is the next step setting this expression equal to 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but Im having problem solving for 0 at this point.

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

and consider the x that will make the equation undefined as a critical number

OpenStudy (anonymous):

make the derivative i mean

OpenStudy (anonymous):

so I have to substitute a value for x that will make this expression ---> (1)/[3(x+2)^(-2/3)]. equal 0?

OpenStudy (anonymous):

your derivative should be 3x^2+12x+12

OpenStudy (anonymous):

if I substitute -2 for x, the entire denominator equals 0. so would -2 b my critical number?

OpenStudy (anonymous):

yes you cant find a number that can make your derivative equal to 0 so -2 is the only critical number

OpenStudy (anonymous):

ohh okay!!! thanks a lot! :)

OpenStudy (anonymous):

@pineda isnt the derivative wrong?

OpenStudy (anonymous):

I think its right. I double checked it on my calculator

OpenStudy (anonymous):

it looks fine for me its \[1/3\sqrt[3]{x+2^{2}}\]

OpenStudy (anonymous):

is the equation \[(x+2)^{3}\]

OpenStudy (anonymous):

pineda100 is right

OpenStudy (anonymous):

its \[\sqrt[3]{x+2}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh oops read it wrong

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

original equation is \[\sqrt[3]{ }x+2\]. sorry i coudlnt get the square root sign to appear completely

OpenStudy (anonymous):

ya then pineda is correct. my bad.

OpenStudy (anonymous):

okay thanks a lot for your help pineda and everyone else who contritbuted! much appreciated! :)